# Exploration of Quantum Computing in Materials Discovery for Direct Air Capture Applications

### v2

@@ 1 Introduction
Technologies for capturing greenhouse gases (GHG) are essential climate mitigation tools that can help reduce GHG atmospheric concentration and limit global warming.1 Existing energy generation infrastructure, such as natural gas and coal-fired power plants, can be retrofitted to perform post-combustion capture of carbon dioxide. This strategy can extend the lifetime of existing infrastructure and minimize the impact of energy generation to the global carbon budget. Nevertheless, merely bringing emissions to zero is not sufficient for achieving climate goals. Carbon dioxide removal from the atmosphere is needed for limiting the temperature increase international2022direct. Direct air capture (DAC) is an emerging carbon removal technology that employs sorbents to capture CO2 directly from the atmosphere.
Liquid sorbent based technologies, such as amine scrubbing, are well-established for capturing carbon dioxide.sanz2016direct Nevertheless, despite their fast adsorption rate and high CO2 saturation, amines have certain disadvantages. They are corrosive, and due to their high desorption energy requirement, are responsible for 60-80% of the total operating costs of CCS.2, sabatino2021comparative Solid sorbent materials for carbon dioxide capture offer an alternative approach to overcome the limitations of amine-based, liquid adsorption. This material class includes a diverse range of nanoporous solids, such as zeolites, metal-organic frameworks, covalent-organic frameworks, zeolitic imidazolate frameworks and porous polymer networks, among others. Porous solid sorbents can selectively adsorb CO2 without bond formation (i.e., via physisorption), which drastically lowers the regeneration energy and, therefore, the operating costs.kumar2015direct On the other hand, solid sorbents can be very sensitive to the presence of water vapor in the atmosphere, leading to decreased selectivity due to competitive adsorption with CO2 and incomplete regeneration due to water passivation of adsorption sites.decoste2013effect
Among the solid materials currently under investigation, Metal-Organic Frameworks (MOFs) have attracted attention due to their customizable porous structure and large chemical diversity.3 MOFs are composed of two types of building blocks, i.e., organic linkers and metal ions, which can be combined to form nanoporous structures with channels whose diameters range from a few to tens of Angstroms. Their chemical and geometrical aspects can be modified, to a large extent, independently for optimizing both the absolute (adsorption capacity) and relative (adsorption selectivity) metrics that determine capture and separation performance.yao2021inverse The large variety of MOFs poses significant computational challenges for the selection of the best candidate material for a given application. There are databases of hypothetical MOFs,4, 5 as well as experimentally validated MOFs,6 containing $10^{5}-10^{6}$ materials in total.
@@
In this work, we perform Quantum Chemistry calculations, on classical and on quantum computing hardware, in the form of the superconductor-based quantum computers available in the IBM Quantum platform, to estimate the binding potential of a set of representative molecules in the atmosphere (CO2, H2O and N2) within a simplified model system. We limited the search space for a carbon capture material by focusing on the X-MOF-74 family (where X=Mg+2, Ni+2, Fe+2, Co+2, Zn+2, Mn+2, Cu+2) as a template for exploring the chemical diversity in MOFs. The MOF-74 family is known for the presence of coordinatively unsaturated metal ions (open metal sites) that can form a bond with adsorbate molecules, greatly influencing the adsorption strength and selectivity, especially at low pressures. The formation of chemical bonds with the adsorbate molecule is a phenomenon known as chemisorption and determines, to a great extent, the desorption energy cost.kumar2015direct, decoste2013effect
@@ 1.1.2 Simulation methods
We chose PySCF,7, 8 as the driver for performing classical benchmarks. We performed two types of classical benchmarks with PySCF: Hartree-Fock (HF) and Density Functional Theory (DFT). The chosen basis set was 6-31g* and was kept consistent across all calculations, including for the quantum calculations. For DFT we chose the B3LYP functional. This particular choice of parameters was motivated by the methodology employed to compute 3D structures (SD-file) of the H2O, CO2, and N2 molecules found in the NIST Chemistry WebBook, from which we also took the molecular structures used in the calculations. 9
@@ 1.1.3 Considerations for Simulations on Quantum Computers
We chose Qiskit10 as the driver for performing quantum simulations using the Variational Quantum Eigensolver (VQE).11, 12, 13 We employed three methodological variants:
  • A standard VQE algorithm with an application-specific, compute-heavy UCCSD ansatz,14 to be treated as the quantum benchmark, as it provides the best possible result for this kind of application. Unfortunately, it is challenging to use this approach on current quantum hardware, due to its complexity.
  • A standard VQE algorithm with a general-purpose, hardware-efficient heuristic ansatz tilly2022variational that can run on current quantum hardware. In our case, we determined that an ansatz with RY and RZ rotations, which we call RYRZ, converges to the intended result despite requiring more iterations.
  • A modified ADAPT-VQE algorithm15 that utilizes qubit operators in the operator pool. This method is called qubit-ADAPT-VQE and has the advantage of giving similar results to UCCSD while having much lower requirements on the quantum hardware.16
@@ 1.1.4 Active Space Selection
To limit the computational cost for the VQE calculations, we performed a suitable active space selection. We used a known method implemented within PySCF called atomic valence active space (AVAS). AVAS is an automated method to construct an active space for electronic structure calculations in strongly correlated chemical systems, where single-determinant wave functions are inadequate. It focuses on identifying specific valence atomic orbitals, such as transition metal d orbitals or those involved in bond dissociation, as a basis for defining the active space. The approach leverages density matrix embedding theory (DMET) to determine the minimal set of orbitals required to capture the essential electronic features, aiming for an active space no more than twice the size of the initially chosen orbitals. This method does not differentiate between “fragment” and “bath” orbitals, instead maintaining the occupied and virtual character of active orbitals, facilitating a natural truncation process to minimize active space size while capturing critical determinants.17
@@ 2.1 Convergence studies
The performance advantage of qubit-ADAPT-VQE comes from the fact that the ansatz grows on demand from an operator pool as the convergence thresholds require. We created the operator pool by first building an UCSSD ansatz and then removing Pauli strings with an even number of Y operations.16 This way we are left only with the “odd” Pauli Strings which are the ones relevant for the energy calculation. We chose the SLSQP optimizer to run the simulations. Finally, we defined the convergence thresholds: maximum number of iterations as $4$, maximum operator gradient as $10^{-4}$, and maximum change of the energy expectation value between consecutive iterations as $10^{-3}$. These convergence thresholds ensure chemical accuracy ($\sim 1\ \text{mHa}$) with respect to the UCCSD-VQE and NumPy results. These settings were used for all qubit-ADAPT-VQE simulations.
@@
We had to choose a suitable active space to run these quantum simulations in a reasonable time. Whenever possible, we utilized the AVAS automated selection method mentioned before, but if the given active space required more than 15 qubits, we would default that calculation to an active space of six electrons and seven molecular orbitals (6,7) chosen around the HOMO-LUMO level with orbitals ordered by energy contribution. This active space requires 10-11 qubits, which we adopted as a reasonable compromise between quality and runtime. All VQE and PySCF simulations were executed on a laptop (Apple MacBook Air M1 2020). We ran the simulations on a slim environment so that anyone with access to our code repository could easily reproduce our results.18
@@ 2.2 Mg+2 systems
The 4-qubit active space limitation led us to focus only on the Mg+2 systems for the VQE calculations on quantum hardware which we refer to as hardware experiments. This active space corresponds to $\approx 5\%$ of the energy contributions to the ground state, therefore we saw no point in simulating systems with a higher electron count as it would mean an even lower representation of the results obtained. We used Qiskit Runtime and ibmq_montreal to run all our quantum hardware experiments, and we employed error mitigation techniques such as Twirled Readout Error Extinction (T-REx).19.
@@
After we calculate the energies, we fit a Morse potential20 to the data points to help us calculate precise equilibrium distances and molecular binding energies. This potential is a suitable inter-atomic model for the energy of diatomic molecules. In the PES plots, we offset the energies so that the asymptotic limit at infinite separation coincides with the zero-energy point to facilitate the extraction of the binding energy.
@@
The experimental binding energies for CO2 on Mg-MOF-74 are in the range of 39-47 kJ/mol.21 These values translate to 15-18 mHa per molecule, which is about 6x smaller than the results we obtained. This illustrates the limits of the oversimplification made by reducing the Mg-MOF-74 structure to a single metal ion.
@@ 3 Code availability
The code used to generate the results presented in this paper is available in a public GitHub repository.18 This repository also includes additional experimental results that were not detailed in the manuscript. For ease of use, the repository contains a README file that explains how the additional results complement those discussed in the paper and guides the navigation of the supplementary material. Should there be any issues or questions regarding the code and its implementation, we encourage readers to reach out.

### v3

@@ 1 Introduction
Technologies for capturing greenhouse gases (GHG) are essential climate mitigation tools that can help reduce GHG atmospheric concentration and limit global warming.1 Existing energy generation infrastructure, such as natural gas and coal-fired power plants, can be retrofitted to perform post-combustion capture of carbon dioxide. This strategy can extend the lifetime of existing infrastructure and minimize the impact of energy generation to the global carbon budget. Nevertheless, merely bringing emissions to zero is not sufficient for achieving climate goals. Carbon dioxide removal from the atmosphere is needed for limiting the temperature increase 2. Direct air capture (DAC) is an emerging carbon removal technology that employs sorbents to capture CO2 directly from the atmosphere.
Liquid sorbent based technologies, such as amine scrubbing, are well-established for capturing carbon dioxide.3 Nevertheless, despite their fast adsorption rate and high CO2 saturation, amines have certain disadvantages. They are corrosive, and due to their high desorption energy requirement, are responsible for 60-80% of the total operating costs of CCS.4, 5 Solid sorbent materials for carbon dioxide capture offer an alternative approach to overcome the limitations of amine-based, liquid adsorption. This material class includes a diverse range of nanoporous solids, such as zeolites, metal-organic frameworks, covalent-organic frameworks, zeolitic imidazolate frameworks and porous polymer networks, among others. Porous solid sorbents can selectively adsorb CO2 without bond formation (i.e., via physisorption), which drastically lowers the regeneration energy and, therefore, the operating costs.6 On the other hand, solid sorbents can be very sensitive to the presence of water vapor in the atmosphere, leading to decreased selectivity due to competitive adsorption with CO2 and incomplete regeneration due to water passivation of adsorption sites.7
Among the solid materials currently under investigation, Metal-Organic Frameworks (MOFs) have attracted attention due to their customizable porous structure and large chemical diversity.8 MOFs are composed of two types of building blocks, i.e., organic linkers and metal ions, which can be combined to form nanoporous structures with channels whose diameters range from a few to tens of Angstroms. Their chemical and geometrical aspects can be modified, to a large extent, independently for optimizing both the absolute (adsorption capacity) and relative (adsorption selectivity) metrics that determine capture and separation performance.9 The large variety of MOFs poses significant computational challenges for the selection of the best candidate material for a given application. There are databases of hypothetical MOFs,10, 11 as well as experimentally validated MOFs,12 containing $10^{5}-10^{6}$ materials in total.
@@
In this work, we perform Quantum Chemistry calculations, on classical and on quantum computing hardware, in the form of the superconductor-based quantum computers available in the IBM Quantum platform, to estimate the binding potential of a set of representative molecules in the atmosphere (CO2, H2O and N2) within a simplified model system. We limited the search space for a carbon capture material by focusing on the X-MOF-74 family (where X=Mg+2, Ni+2, Fe+2, Co+2, Zn+2, Mn+2, Cu+2) as a template for exploring the chemical diversity in MOFs. The MOF-74 family is known for the presence of coordinatively unsaturated metal ions (open metal sites) that can form a bond with adsorbate molecules, greatly influencing the adsorption strength and selectivity, especially at low pressures. The formation of chemical bonds with the adsorbate molecule is a phenomenon known as chemisorption and determines, to a great extent, the desorption energy cost.6, 7
@@ 1.1.2 Simulation methods
We chose PySCF,13, 14 as the driver for performing classical benchmarks. We performed two types of classical benchmarks with PySCF: Hartree-Fock (HF) and Density Functional Theory (DFT). The chosen basis set was 6-31g* and was kept consistent across all calculations, including for the quantum calculations. For DFT we chose the B3LYP functional. This particular choice of parameters was motivated by the methodology employed to compute 3D structures (SD-file) of the H2O, CO2, and N2 molecules found in the NIST Chemistry WebBook, from which we also took the molecular structures used in the calculations. 15
@@ 1.1.3 Considerations for Simulations on Quantum Computers
We chose Qiskit16 as the driver for performing quantum simulations using the Variational Quantum Eigensolver (VQE).17, 18, 19 We employed three methodological variants:
  • A standard VQE algorithm with an application-specific, compute-heavy UCCSD ansatz,20 to be treated as the quantum benchmark, as it provides the best possible result for this kind of application. Unfortunately, it is challenging to use this approach on current quantum hardware, due to its complexity.
  • A standard VQE algorithm with a general-purpose, hardware-efficient heuristic ansatz 21 that can run on current quantum hardware. In our case, we determined that an ansatz with RY and RZ rotations, which we call RYRZ, converges to the intended result despite requiring more iterations.
  • A modified ADAPT-VQE algorithm22 that utilizes qubit operators in the operator pool. This method is called qubit-ADAPT-VQE and has the advantage of giving similar results to UCCSD while having much lower requirements on the quantum hardware.23
@@ 1.1.4 Active Space Selection
To limit the computational cost for the VQE calculations, we performed a suitable active space selection. We used a known method implemented within PySCF called atomic valence active space (AVAS). AVAS is an automated method to construct an active space for electronic structure calculations in strongly correlated chemical systems, where single-determinant wave functions are inadequate. It focuses on identifying specific valence atomic orbitals, such as transition metal d orbitals or those involved in bond dissociation, as a basis for defining the active space. The approach leverages density matrix embedding theory (DMET) to determine the minimal set of orbitals required to capture the essential electronic features, aiming for an active space no more than twice the size of the initially chosen orbitals. This method does not differentiate between “fragment” and “bath” orbitals, instead maintaining the occupied and virtual character of active orbitals, facilitating a natural truncation process to minimize active space size while capturing critical determinants.24
@@ 2.1 Convergence studies
The performance advantage of qubit-ADAPT-VQE comes from the fact that the ansatz grows on demand from an operator pool as the convergence thresholds require. We created the operator pool by first building an UCSSD ansatz and then removing Pauli strings with an even number of Y operations.23 This way we are left only with the “odd” Pauli Strings which are the ones relevant for the energy calculation. We chose the SLSQP optimizer to run the simulations. Finally, we defined the convergence thresholds: maximum number of iterations as $4$, maximum operator gradient as $10^{-4}$, and maximum change of the energy expectation value between consecutive iterations as $10^{-3}$. These convergence thresholds ensure chemical accuracy ($\sim 1\ \text{mHa}$) with respect to the UCCSD-VQE and NumPy results. These settings were used for all qubit-ADAPT-VQE simulations.
@@
We had to choose a suitable active space to run these quantum simulations in a reasonable time. Whenever possible, we utilized the AVAS automated selection method mentioned before, but if the given active space required more than 15 qubits, we would default that calculation to an active space of six electrons and seven molecular orbitals (6,7) chosen around the HOMO-LUMO level with orbitals ordered by energy contribution. This active space requires 10-11 qubits, which we adopted as a reasonable compromise between quality and runtime. All VQE and PySCF simulations were executed on a laptop (Apple MacBook Air M1 2020). We ran the simulations on a slim environment so that anyone with access to our code repository could easily reproduce our results.25
@@ 2.2 Mg+2 systems
The 4-qubit active space limitation led us to focus only on the Mg+2 systems for the VQE calculations on quantum hardware which we refer to as hardware experiments. This active space corresponds to $\approx 5\%$ of the energy contributions to the ground state, therefore we saw no point in simulating systems with a higher electron count as it would mean an even lower representation of the results obtained. We used Qiskit Runtime and ibmq_montreal to run all our quantum hardware experiments, and we employed error mitigation techniques such as Twirled Readout Error Extinction (T-REx).26.
@@
After we calculate the energies, we fit a Morse potential27 to the data points to help us calculate precise equilibrium distances and molecular binding energies. This potential is a suitable inter-atomic model for the energy of diatomic molecules. In the PES plots, we offset the energies so that the asymptotic limit at infinite separation coincides with the zero-energy point to facilitate the extraction of the binding energy.
@@
The experimental binding energies for CO2 on Mg-MOF-74 are in the range of 39-47 kJ/mol.28 These values translate to 15-18 mHa per molecule, which is about 6x smaller than the results we obtained. This illustrates the limits of the oversimplification made by reducing the Mg-MOF-74 structure to a single metal ion.
@@ 3 Code availability
The code used to generate the results presented in this paper is available in a public GitHub repository.25 This repository also includes additional experimental results that were not detailed in the manuscript. For ease of use, the repository contains a README file that explains how the additional results complement those discussed in the paper and guides the navigation of the supplementary material. Should there be any issues or questions regarding the code and its implementation, we encourage readers to reach out.